OpenStudy (anonymous):

What is the limit of f(x) = x +1/11 − x^2 when x approaches 3? 0.5 2 3 A limit does not exist.

OpenStudy (anonymous):

Replace x for 3

OpenStudy (anonymous):

is it 0.5

OpenStudy (anonymous):

or is it 9

OpenStudy (anonymous):

@kirbykirby hey, could you help this fellow over here? Thanks. I working with someone else at the moment.

OpenStudy (kirbykirby):

\[\frac{x+1}{11-x^2} \] I'm assuming this is the given expression?

OpenStudy (anonymous):

yes

OpenStudy (kirbykirby):

So if plugging in x=3 directly gives you a number and nothing tricky (like divided by 0, or 0/0, or infinity/infinity, etc.), then that's all there is to it :) \[\large \lim_{x\rightarrow 3} \frac{x+1}{11-x^2} =\frac{3+1}{11-3^2}\]

OpenStudy (anonymous):

thank you for the help

OpenStudy (kirbykirby):

=]